# Trajectory Optimization with Global Yaw Parameterization for Field-of-View Constrained Autonomous Flight

### v1

@@ I Introduction
The representation and parameterization of a robot’s state play crucial roles in motion planning and control. While robot positions can be uniquely formulated in global coordinates, available representations for attitude can be problematic because of singularity and continuity issues. The Euler angles vectors, quaternions, and rotation matrices are widely used in control within the local domain of definition but may introduce continuity or stability issues for global state space [12, 13]. High-dimensional rotation representations [14] or methods [15] that use multiple charts to cover the $SO(3)$ manifold provide global parameterization; nevertheless, they significantly complicate the tasks of encoding perception-based constraints, as well as minimizing the control effort in trajectory optimization. The exponential map [16] can achieve almost global parameterization and can also be used for trajectory generation. However, it needs an additional selection of the range of the angle bounds for the optimization initialization. The quality of the optimization is directly affected by such selections. In practice, finding a suitable selection for a particular application scenario is challenging.
Figure 1: Snapshots of hardware experiment of traversal planning with yaw constraints. We use our SWaP-constrained Falcon250 quadrotor platform and Scarab ground robot platform as a tracking target (V-D). The video is available at: https://youtu.be/TtgMTXtyhkE
@@ II-B Yaw Trajectory Planning
Leveraging the property of differential flatness, the complete dynamics of the quadrotor can be efficiently represented through flat outputs, and position and yaw trajectories can be parameterized by piecewise polynomials that can be optimized[17]. Perception-related constraints or objectives usually couple all flat outputs, introducing nonlinearity and modeling difficulties in the problem formulation. As position and yaw trajectories can be generated independently for tracking in the output space, separate yaw optimization with approximated perception constraints is extensively applied in many works. Some previous works set the trajectory of yaw angles to values determined by a heuristic method, without optimizing it further [6, 7, 8, 19, 9, 25].
@@ III-B Differential-flatness-based Trajectory Optimization
(Differential Flatness for Control-affine Systems[28]) A dynamical system of the form $\dot{x}=f(x)+g(x)u$ is said to have a flatness mapping if $x,u$ can be represented by the “flat outputs” and a finite number ($\kappa$) of their derivatives, $(\mathbf{x}(t),\mathbf{u}(t))=\Psi(\bm{\sigma}^{[\kappa]}(t))$, where $\bm{\sigma}^{[\kappa]}(t)=(\bm{\sigma}(t),\dot{\bm{\sigma}}(t),\cdots,\bm{\sigma}^{(\kappa-1)}(t))$.
@@ III-C Perception-based Trajectory Generation
$\bm{\sigma}^{*}=\operatorname*{arg\,min}\{c(\bm{\sigma})|\ \bm{\sigma}(t_{0})=x_{s},\gamma(\bm{\sigma}(t_{i}),w_{i})\leq 0,$ (3)
@@ IV-A Parameterization with Direct Mapping
Directly solving the initial problem (LABEL:eq:optimization) would result in nonlinear constraints and objective functions. We will further discuss both convex optimization formulation via approximation for yaw trajectories and joint nonlinear optimization to efficiently generate coupled trajectories.
@@ IV-B Optimization with Virtual Variables
Given initial and terminal states, with a sequence of intermediate heading vectors $[\bm{\psi}_{1},...,\bm{\psi}_{M-1}]^{T}$, the minimum-control yaw trajectory generation problem can be stated as
@@ IV-C Traversal Planning via Nonlinear Optimization
$\min_{\bm{\sigma}_{v}(t)}J_{c}(\bm{\sigma}_{v}(t))+\sum I(\mathcal{G}(\bm{\sigma}_{v}(t),\cdot)),$ (12)
where $J_{c}(\cdot)$ is the total control effort. The relaxation of inequality constraints can be achieved through the method presented in [25], applying a time integral penalty functional denoted as $I(\cdot)$. We follow similar cost formulations for position-only constraints and focus on the discussion of perception-related penalties in the following context. For cost evaluation, we analyze the original yaw dynamics,
$\dot{\bm{\psi}}(t)=\left({\rm I}_{2}-\frac{\bm{s}(t)\bm{s}(t)^{T}}{||\bm{s}(t)||^{2}}\right)\frac{\bm{s}(t)}{||\bm{s}(t)||},$ (13)
@@ V-C A Note on an Alternative Representation
$\langle\begin{bmatrix}\cos(\psi(t_{i}))\\ \sin(\psi(t_{i}))\end{bmatrix},\begin{bmatrix}\cos(\psi_{i}))\\ \sin(\psi_{i}))\end{bmatrix}\rangle=1,\forall i\in{1,...,M},$ (20d)

### v2

@@ I Introduction
The representation and parameterization of a robot’s state play crucial roles in motion planning and control. While robot positions can be uniquely formulated in global coordinates, available representations for attitude can be problematic because of singularity and continuity issues. The Euler angles vectors, quaternions, and rotation matrices are widely used in control within the local domain of definition but may introduce continuity or stability issues for global state space [12, 13]. High-dimensional rotation representations [14] or methods [15] that use multiple charts to cover the $SO(3)$ manifold provide global parameterization; nevertheless, they significantly complicate the tasks of encoding perception-based constraints, as well as minimizing the control effort in trajectory optimization. The exponential map [16] provides an almost global parameterization which can also be utilized for trajectory generation. However, it needs an additional selection of the range of the angle bounds for the optimization initialization. The quality of the optimization is directly affected by such selections. In practice, finding a suitable selection for a particular application scenario is challenging.
Figure 1: Snapshots of hardware experiment of traversal planning with yaw constraints. We use our SWaP-constrained Falcon250 quadrotor platform and Scarab ground robot platform as a tracking target (V-D). The video is available at: https://youtu.be/QMbkoTyKe-k
@@ II-B Yaw Trajectory Planning
Leveraging the property of differential flatness, the complete dynamics of the quadrotor can be efficiently represented through flat outputs, and position and yaw trajectories can be parameterized by piecewise polynomials that can be optimized[17]. Perception-related constraints or objectives usually couple all flat outputs, introducing nonlinearity and modeling difficulties in the problem formulation. As position and yaw trajectories can be generated independently with certain dynamic constraints for tracking in the output space, separate yaw optimization with approximated perception constraints is extensively applied in many works. Some previous works set the trajectory of yaw angles to values determined by a heuristic method, without optimizing it further [6, 7, 8, 19, 9, 25].
@@ III-B Differential-flatness-based Trajectory Optimization
(Differential Flatness for Control-affine Systems[28]) A dynamical system of the form $\dot{x}=f(x)+g(x)u$ is said to have a flatness mapping if $x,u$ can be represented by the “flat outputs” and a finite number ($\kappa$) of their derivatives, $(\mathbf{x}(t),\mathbf{u}(t))=\Psi(\bm{\sigma}^{[\kappa]}(t))$, where $\bm{\sigma}^{[\kappa]}(t)=(\bm{\sigma}(t),\dot{\bm{\sigma}}(t),\cdots,\bm{\sigma}^{(\kappa)}(t))$.
@@ III-C Perception-based Trajectory Generation
$\bm{\sigma}^{*}=\operatorname*{arg\,min}\{c(\bm{\sigma})|\ \bm{\sigma}(t_{0})=x_{s},\gamma(\bm{\sigma}(t_{i}),\bm{w_{i}})\leq 0,$ (3)
@@ IV-A Parameterization with Direct Mapping
Directly solving the initial problem (LABEL:eq:_yaw_opt) would result in nonlinear constraints and objective functions. We will further discuss both convex optimization formulation via approximation for yaw trajectories and joint nonlinear optimization to efficiently generate coupled trajectories.
@@ IV-B Optimization with Virtual Variables
Given initial and terminal states, with a sequence of intermediate heading vectors $[\bm{\psi}_{1},...,\bm{\psi}_{M-1}]^{T}$, the minimum-yaw-control trajectory generation problem with dynamic bounds can be stated as
@@ IV-C Traversal Planning via Nonlinear Optimization
$\min_{\bm{\sigma}_{v}(t)}J_{c}(\bm{\sigma}_{v}(t))+\sum_{i}\lambda_{i}I(\mathcal{G}_{i}(\bm{\sigma}_{v}(t),\cdot)),$ (12)
where $J_{c}(\cdot)$ is the total control effort, and $\mathcal{G}_{i}$ represent different inequality constraints, each weighted by $\lambda_{i}$. The relaxation of inequality constraints can be achieved through the method presented in [25], applying a time integral penalty functional denoted as $I(\cdot)$. We follow similar cost formulations for position-only constraints and focus on the discussion of perception-related penalties in the following context. For cost evaluation, we analyze the original yaw dynamics,
$\dot{\bm{\psi}}(t)=\left({\rm I}_{2}-\frac{\bm{s}(t)\bm{s}(t)^{T}}{||\bm{s}(t)||^{2}}\right)\frac{\dot{\bm{s}}(t)}{||\bm{s}(t)||},$ (13)
@@ V-C A Note on an Alternative Representation
$\begin{pmatrix}\cos(\psi(t_{i}))\\ \sin(\psi(t_{i}))\end{pmatrix}\cdot\begin{pmatrix}\cos(\psi_{i}))\\ \sin(\psi_{i}))\end{pmatrix}=1,\forall i\in{1,...,M},$ (20d)